OpenStudy (anonymous):

Tricky Probability Question: A bag consists of 1 red marble and 4 green marble. If two marbles are drawn successively without replacement, what is the probability of getting red one??

OpenStudy (anonymous):

why is this tricky?

OpenStudy (anonymous):

because two draws one marble

OpenStudy (anonymous):

\[1-\frac{4}{5}\times \frac{3}{4}\]

OpenStudy (anonymous):

8/20 = 2/5

OpenStudy (anonymous):

i get \(\frac 2 5\)

OpenStudy (anonymous):

you solved by calculating the probability of getting green marble

OpenStudy (anonymous):

i solved it by calculating the probability of not getting the red one and subtracting the result from 1

OpenStudy (anonymous):

tricky answer ... than you

OpenStudy (anonymous):

you could also compute \(\frac{1}{5}\times \frac{4}{4}+\frac{4}{5}\times \frac{1}{4}=\frac{1}{5}+\frac{1}{5}=\frac{2}{5}\)

OpenStudy (anonymous):

ya that also is